### v1

# EVDA: Evolving Deepfake Audio Detection Continual Learning Benchmark

Abstract
The rise of advanced large language models such as GPT-4, GPT-4o, and the Claude family has made fake audio detection increasingly challenging. Traditional fine-tuning methods struggle to keep pace with the evolving landscape of synthetic speech, necessitating continual learning approaches that can adapt to new audio while retaining the ability to detect older types. Continual learning, which acts as an effective tool for detecting newly emerged deepfake audio while maintaining performance on older types, lacks a well-constructed and user-friendly evaluation framework. To address this gap, we introduce EVDA, a benchmark for evaluating continual learning methods in deepfake audio detection. EVDA includes classic datasets from the Anti-Spoofing Voice series, Chinese fake audio detection series, and newly generated deepfake audio from models like GPT-4 and GPT-4o. It supports various continual learning techniques, such as Elastic Weight Consolidation (EWC), Learning without Forgetting (LwF), and recent methods like Regularized Adaptive Weight Modification (RAWM) and Radian Weight Modification (RWM). Additionally, EVDA facilitates the development of robust algorithms by providing an open interface for integrating new continual learning methods.
1 Introduction
Fake audio detection has become increasingly crucial in recent years due to the rapid development of deep learning models, particularly large language models such as GPT-4 (OpenAI, 2023), GPT-4o, and Claude family (Anthropic, 2024). These models have significantly blurred the line between real and generated speech, making it challenging for human ears to distinguish between them. With the accessibility of tools like GPT-4o to the general public, generating synthetic audio has become effortless, presenting significant challenges for societal security and safety.
The continuous development of large language models like GPT-4, GPT-4o, and others has led to the emergence of new techniques for generating synthetic speech. This evolution has made it increasingly difficult for humans to discern between generated and authentic speech. The availability of GPT-4o to free users has further democratized the creation of fake audio, presenting both convenience and significant challenges for various domains such as security.
Continual learning, particularly in the context of fake audio detection, has emerged as an effective approach to address the evolving landscape of synthetic speech generation. By continuously collecting data from newly emerging techniques, including those generated by large language models, continual learning enables models to adapt to new types of synthetic speech while retaining the ability to detect previously encountered ones.
For example, traditional model training and fine-tuning techniques may result in the forgetting of older models’ characteristics once the model learns to distinguish newer models like GPT-4o. In contrast, continual learning ensures that the model retains the ability to discern older models like GPT-4 while learning to identify newer ones.
To address the challenges posed by evolving large models and speech generation techniques, we propose a continually updated benchmark for EVoloving synthetic and Deepfake Audio detection using various continual learning methods, named EVDA. Our benchmark includes classic datasets from the Anti-Spoofing Voice series challenges (Wu et al., 2015; Todisco et al., 2019; Liu et al., 2023) for detection tasks, as well as Chinese fake audio detection datasets such as Audio Deepfake Detection series (ADD 2022, ADD 2023)(Yi et al., 2022; 2023), among others and newly emerging deepfake audio generated by recent large language models, such GPT-4 and GPT-4o.
Additionally, we plan to incorporate fake audio generated by models like GPT-4, GPT-4o, Claude, and future large models as new detection tasks. On the continual learning front, we support classic methods like Elastic Weight Consolidation (EWC)(Kirkpatrick et al., 2017), Learning without Forgetting (LwF)(Li & Hoiem, 2017), as well as recent methods like Regularized Adaptive Weight Modification (RAWM)(Zhang et al., 2023), Radian Weight Modification (RWM)(Zhang et al., 2024). Furthermore, we provide an open interface to allow researchers to integrate their own continual learning methods into the benchmark, facilitating the development of more effective and robust algorithms to tackle the continuous evolution of fake audio generated by large models.
The primary contribution of this paper is EVDA, a benchmark designed to evaluate continual learning methods in the context of evolving deepfake audio detection. EVDA encompasses both classic and newly emerging datasets, and supports various continual learning algorithms. We emphasize the importance of continual learning as a solution to address the dynamic nature of these challenges.
2 Proposed Method
2.1 Data Architecture
Our benchmark architecture consists of twelve experiences (Exp_1 to Exp_12), each representing different datasets and conditions for detecting synthetic audio. These experiences vary in data sources, language, acoustic conditions, and the number of training and evaluation samples. The data architecture of EVDA benchmark are listed as Table 1.
Table 1: The Data architecture of the Evda deepfake audio detection benchmark. T: True; F: False
Task $\mathbf{Task_{1}}$ $\mathbf{Task_{2}}$ $\mathbf{Task_{3}}$ $\mathbf{Task_{4}}$
Source ASVspoof2015 ASVspoof2019 FoR HAD
Sub Source - LA - -
Have Vocoder ID T T F F
Language English English English Chinese
Acoustic Condition No significant noise Logical access Human Human/Partially fake
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
Task $\mathbf{Task_{5}}$ $\mathbf{Task_{6}}$ $\mathbf{Task_{7}}$ $\mathbf{Task_{8}}$
Source ASVspoof2021 WaveFake FMFCC-A ADD 2022
Sub Source LA - - -
Have Vocoder ID T T F F
Language English English/Japanese Chinese Chinese
Acoustic Condition Logical access Human Human Low-quality/Partially fake
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
Task $\mathbf{Task_{9}}$ $\mathbf{Task_{1}0}$ $\mathbf{Task_{11}}$ $\mathbf{Task_{12}}$
Source In-the-Wild ADD 2023 GPT-4 GPT-4o
Sub Source - - - -
Have Vocoder ID F F F F
Language English Chinese English English
Acoustic Condition Real world Low-quality/Partially fake Logical access Logical access
Training Samples 2000 2000 2000 2000
Testing Samples 500 500 500 500
The proposed benchmark architecture, as outlined in the table, is designed to comprehensively evaluate the performance of audio spoofing and anti-spoofing models across a diverse set of conditions and datasets. Each experiment (Exp_1 to Exp_12) has been carefully crafted to test various aspects of the detection models under different acoustic and linguistic environments, source integrity, and data authenticity. Below, we detail each experiment’s setup and objectives:
  • Exp_1: ASVspoof2015 (English)(Wu et al., 2015) - Utilizes 2000 training and 500 evaluation samples from the ASVspoof2015 dataset, focusing on English language audio in clean conditions without significant channel or background noise effects. Vocoder ID is available, aiding in more precise synthetic speech detection.
  • Exp_2: ASVspoof2019 - Logical Access (LA) (English)(Todisco et al., 2019) - Also incorporates 2000 training and 500 evaluation samples but focuses on logical access scenarios to test models against sophisticated spoofing attacks in English. Vocoder ID is again available, providing critical data for assessing synthetic signal traits.
  • Exp_3 and Exp_4: FoR (Reimao & Tzerpos, 2019) and HAD (Yi et al., 2021) (English and Chinese) - Each uses 2000 training and 500 evaluation samples. FoR involves purely human-generated audio in English, while HAD includes a mix of human and partially fake audio in Chinese, presenting challenges in cross-lingual and authenticity discernment.
  • Exp_5: ASVspoof2021 - Logical Access (English)(Liu et al., 2023) - Continues the logical access examination with the latest dataset iteration, using English audio samples. The presence of vocoder ID helps in identifying new spoofing techniques that have emerged since the 2019 dataset.
  • Exp_6: WaveFake (English/Japanese)(Frank & Schönherr, 2021) - Addresses bilingual detection scenarios, testing the robustness of models against spoofs in both English and Japanese. This dataset’s inclusion highlights the benchmark’s focus on multilingual and cross-cultural spoofing detection.
  • Exp_7 and Exp_8: FMFCC-A (Zhang et al., 2021) and ADD 2022 (Yi et al., 2022) (Chinese) - Each experiment uses 2000 training and 500 evaluation samples of Chinese audio. These tests are critical for models to adapt to low-quality, partially fake audio environments and a generation & detection game setting, increasing the difficulty of the spoofing detection task.
  • Exp_9 and Exp_10: In-the-Wild (Müller et al., 2022) and ADD 2023 (Yi et al., 2023) (English and Chinese) - These experiments push the boundaries by including real-world audio scenarios (Exp_9) and the latest generation & detection games in a low-quality, partially fake setting (Exp_10), testing the models’ effectiveness in dynamically changing and challenging acoustic environments.
  • Exp_11: Collected Audio Generated by GPT-4 (English) (OpenAI, 2023) - This experiment involves 2000 training and 500 evaluation samples generated by the GPT-4 model, tailored to produce synthetic speech that mimics human conversational patterns in English. The experiment aims to evaluate the effectiveness of detection models against the latest AI-generated audio, challenging them to distinguish between human-like synthetic speech and authentic human speech. The sophisticated natural language generation capabilities of GPT-4 represent a significant leap in the realism of generated audio, thus presenting a unique challenge for anti-spoofing technologies.
  • Exp_12: Collected Audio Generated by GPT-4o (English) - Similar to Exp_11, this experiment also uses 2000 training and 500 evaluation samples, but the audio is generated by GPT-4o, an optimized version of GPT-4 designed to handle English languages with enhanced contextual understanding and expression. This experiment tests the models’ abilities to detect synthetic speech generated by GPT-4o, thereby assessing their adaptability and robustness in a globally interconnected and linguistically diverse world.
The benchmark is designed not only to assess the current state-of-the-art in spoofing detection but also to push the development of more robust, adaptable, and cross-lingual models capable of handling the complexities of real-world audio spoofing scenarios.
2.2 Data Selection Methodology
For each task within our benchmark, we employ two strategies for data selection:
  1. 1.
    Random Selection: We randomly select a fixed number of authentic and spoofed audio samples from the original datasets or collected data. Initially, we split the data into two groups: Real and Fake. From the ”Real” group, we randomly select 1000 real samples, and from the ”Fake” group, we select 1000 fake samples corresponding to various generation methods. These samples are combined to form the training set. For the evaluation set, we follow the same procedure but select only 250 real and 250 fake samples (1/4 of the training set size). This approach ensures a balanced representation of real and fake samples in both the training and evaluation sets.
  2. 2.
    Most Informative Selection: In this method, we aim to select the most informative samples from the original datasets. First, we train $n$ discriminative models for spoof detection using external datasets, ensuring these models achieve extremely low loss on their respective training sets. These well-trained models act as experts during the sample selection process. Each sample from the original dataset is evaluated by all experts, and the samples are scored based on the entropy of the experts’ predictions using the formula:
    $\text{Entropy}=-(p_{1}\log(p_{1})+p_{2}\log(p_{2}))$ (1)
    where
    $p_{1}=\frac{N_{\text{real}}}{N_{\text{real}}+N_{\text{fake}}},\quad p_{2}=\frac{N_{\text{fake}}}{N_{\text{real}}+N_{\text{fake}}}$ (2)
    Here, $N_{\text{real}}$ and $N_{\text{fake}}$ represent the number of models classifying the sample as real or fake, respectively. According to information theory (Shannon, 1948), samples with the highest entropy values contain the most information, indicating uncertainty and hence, difficulty in classification. We list the entropy of each sample and rank them from highest to lowest. The top 250 real and 250 fake samples are selected to construct the evaluation set, ensuring it includes the most challenging and informative samples. The next 1000 real and 1000 fake samples are selected for the training set, which benefits from the inclusion of diverse and informative samples.
2.3 Continuous Learning Methods in Benchmark
The benchmark incorporates a variety of continuous learning approaches to evaluate and enhance the adaptability of audio spoofing detection systems. These methods include:
  • CWRStar(Maltoni & Lomonaco, 2019), Replay(Rolnick et al., 2019), Joint Training(Parisi et al., 2019): Basic strategies for integrating new knowledge while retaining previously learned information without severe forgetting.
  • GDumb (Finetuning)(Prabhu et al., 2020), (Bias)Reservoir(Vitter, 1985), Cumulative: Techniques focusing on selectively updating the model or maintaining a subset of previous data to balance between old and new learning.
  • LwF(Li & Hoiem, 2017), OWM(Zeng et al., 2019), RAWM(Zhang et al., 2023), RWM(Zhang et al., 2024), GEM(Lopez-Paz & Ranzato, 2017), AGEM(Chaudhry et al., 2019): Methods that leverage gradients or error restrictions to mitigate catastrophic forgetting.
  • EWC(Kirkpatrick et al., 2017), Finetune, ICaRL(Rebuffi et al., 2017), SIZenke et al. (2017), AR1, StreamingLDA: Advanced strategies that introduce regularization, importance sampling, or dynamically adjust model architecture based on incoming data streams.
In summary, our benchmark is designed to rigorously assess the effectiveness of various continuous learning methods in the context of audio spoofing detection. By incorporating a diverse set of strategies, ranging from basic integration techniques to advanced regularization and dynamic adjustment methods, we aim to provide a comprehensive evaluation framework. This approach not only facilitates the identification of strengths and weaknesses of each method but also paves the way for future advancements in the continual learning domain. The ultimate goal is to develop robust and adaptive models that can efficiently handle the evolving nature of audio spoofing attacks, ensuring high reliability and accuracy in real-world applications.
3 Result
We evaluate our benchmark using the Equal Error Rate (EER) (Wu et al., 2015) metric on seven tasks using the LwF method. All methods and tasks are trained using only-1 linear layer model. The EER results are illustrated as follows:
EER Task1 Task2 Task3 Task4 Task5 Task6 Task7
Train after Task1 0.26 0.334 0.4027 0.428 0.428 0.4847 0.4857
Train after Task2 0.478 $\uparrow$ 0.4836 $\uparrow$ 0.4844 $\uparrow$ 0.4887 $\uparrow$ 0.4893 $\uparrow$ 0.4972 $\uparrow$ 0.522 $\uparrow$
Train after Task3 0.5539 $\uparrow$ 0.5818 $\uparrow$ 0.6064 $\uparrow$ 0.6282 $\uparrow$ 0.6478 $\uparrow$ 0.6654 $\uparrow$ 0.6813 $\uparrow$
Train after Task4 0.6504 $\uparrow$ 0.6221 $\uparrow$ 0.5962 $\downarrow$ 0.5723 $\downarrow$ 0.5503 $\downarrow$ 0.5299 $\downarrow$ 0.511 $\downarrow$
Train after Task5 0.4988 $\downarrow$ 0.496 $\downarrow$ 0.4972 $\downarrow$ 0.497 $\downarrow$ 0.4968 $\downarrow$ 0.4973 $\downarrow$ 0.4939 $\downarrow$
Train after Task6 0.495 $\downarrow$ 0.4949 $\downarrow$ 0.4949 $\downarrow$ 0.4956 $\downarrow$ 0.4959 $\downarrow$ 0.498 $\uparrow$ 0.497 $\uparrow$
Train after Task7 0.4922 $\downarrow$ 0.4915 $\downarrow$ 0.4914 $\downarrow$ 0.4915 $\downarrow$ 0.4909 $\downarrow$ 0.4972 $\downarrow$ 0.4984 $\uparrow$
Table 2: Training Performance After Each Task
Table 2 summarizes the training performance of the model after each task in a sequential learning scenario. The evaluation metric used is Equal Error Rate (EER), and tasks are denoted from Task1 to Task7. The table reveals the model’s performance in terms of EER after training on each task sequentially. Results show that initially, after Task1, the EER is 0.26, gradually improving to 0.4857 after Task7. Notably, training after Task2 leads to a significant improvement, marked with an upward arrow, and the trend continues until Task4. However, after Task4, there is a decline in performance (marked with downward arrows) which persists through Task7. This suggests that the model might start to overfit or lose generalization capabilities after learning from Task4 onwards. Additionally, Task6 shows a slight improvement, suggesting some recovery in performance. Overall, the table provides insights into how the model’s performance evolves over sequential tasks, indicating both improvements and potential challenges in continual learning scenarios.
(a) Task 1
(b) Task 2
(c) Task 3
(d) Task 4
(e) Task 5
(f) Task 6
(g) Task 7
Figure 1: Evaluation on Tasks
4 Conclusion
In conclusion, the rapid evolution of audio generation models poses significant challenges for fake audio detection. However, by adopting continual learning approaches and establishing a comprehensive benchmark, we can effectively address these challenges. Our proposed benchmark provides a standardized platform for evaluating detection methods against evolving synthetic audio techniques, facilitating the development of more robust and effective detection algorithms.