### v3

# Towards Robust Audio Deepfake Detection: A Evolving Benchmark for Continual Learning

Abstract
The rise of advanced large language models such as GPT-4, GPT-4o, and the Claude family has made fake audio detection increasingly challenging. Traditional fine-tuning methods struggle to keep pace with the evolving landscape of synthetic speech, necessitating continual learning approaches that can adapt to new audio while retaining the ability to detect older types. Continual learning, which acts as an effective tool for detecting newly emerged deepfake audio while maintaining performance on older types, lacks a well-constructed and user-friendly evaluation framework. To address this gap, we introduce EVDA, a benchmark for evaluating continual learning methods in deepfake audio detection. EVDA includes classic datasets from the Anti-Spoofing Voice series, Chinese fake audio detection series. It supports various continual learning techniques, such as Elastic Weight Consolidation (EWC), Learning without Forgetting (LwF), and recent continual learning methods for audio deepfake detection like Regularized Adaptive Weight Modification (RAWM) and Radian Weight Modification (RWM). Additionally, EVDA facilitates the development of robust algorithms by providing an open interface for integrating new continual learning methods. Our code will be available at https://github.com/Cecile-hi/Evolving-FAD-CL-Benchmark.git.
1 Introduction
Fake audio detection has become increasingly crucial in recent years due to the rapid development of deep learning models, particularly large language models such as GPT-4 (OpenAI, 2023), GPT-4o, and Claude family (Anthropic, 2024). These models have significantly blurred the line between real and generated speech, making it challenging for human ears to distinguish between them. With the accessibility of tools like GPT-4o to the general public, generating synthetic audio has become effortless, presenting significant challenges for societal security and safety.
The continuous development of large language models like GPT-4, GPT-4o, and others has led to the emergence of new techniques for generating synthetic speech. This evolution has made it increasingly difficult for humans to discern between generated and authentic speech. The availability of GPT-4o to free users has further democratized the creation of fake audio, presenting both convenience and significant challenges for various domains such as security.
Continual learning, particularly in the context of fake audio detection, has emerged as an effective approach to address the evolving landscape of synthetic speech generation. By continuously collecting data from newly emerging techniques, including those generated by large language models, continual learning enables models to adapt to new types of synthetic speech while retaining the ability to detect previously encountered ones.
For example, traditional model training and fine-tuning techniques may result in the forgetting of older models’ characteristics once the model learns to distinguish newer models like GPT-4o. In contrast, continual learning ensures that the model retains the ability to discern older models like GPT-4 while learning to identify newer ones.
To address the challenges posed by evolving large models and speech generation techniques, we propose a continually updated benchmark for EVoloving synthetic and Deepfake Audio detection using various continual learning methods, named EVDA. Our benchmark includes classic datasets from the Anti-Spoofing Voice series challenges (Wu et al., 2015; Todisco et al., 2019; Liu et al., 2023) for detection tasks, as well as Chinese fake audio detection datasets such as Audio Deepfake Detection series (ADD 2022, ADD 2023)(Yi et al., 2022; 2023), among others and newly emerging deepfake audio generated by recent large language models, such GPT-4 and GPT-4o.
Additionally, we plan to incorporate fake audio generated by models like GPT-4, GPT-4o, Claude, and future large models as new detection tasks. On the continual learning front, we support classic methods like Elastic Weight Consolidation (EWC)(Kirkpatrick et al., 2017), Learning without Forgetting (LwF)(Li & Hoiem, 2017), as well as recent methods like Regularized Adaptive Weight Modification (RAWM)(Zhang et al., 2023), Radian Weight Modification (RWM)(Zhang et al., 2024). Furthermore, we provide an open interface to allow researchers to integrate their own continual learning methods into the benchmark, facilitating the development of more effective and robust algorithms to tackle the continuous evolution of fake audio generated by large models.
The primary contribution of this paper is EVDA, a benchmark designed to evaluate continual learning methods in the context of evolving deepfake audio detection. EVDA encompasses both classic and newly emerging datasets, and supports various continual learning algorithms. We emphasize the importance of continual learning as a solution to address the dynamic nature of these challenges.
2 Related Work
2.1 Audio Deepfake Detection
The rise of synthetic audio, often referred to as deepfake audio, has led to significant advancements in detection techniques aimed at distinguishing between genuine and manipulated audio content. Early methods primarily focused on spectral analysis and traditional machine learning approaches, leveraging features such as Mel-frequency cepstral coefficients (MFCCs) and Gaussian Mixture Models (GMMs) to detect anomalies indicative of synthetic audio (Wu et al., 2015). With the evolution of deep learning, more sophisticated models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), have been developed to capture subtle temporal and spectral differences between real and fake audio. Recent advancements have introduced hybrid models that combine deep learning with signal processing techniques to improve detection accuracy in challenging environments, such as noisy or low-quality recordings (Todisco et al., 2019; Liu et al., 2023). Additionally, cross-lingual and cross-dataset generalization has become a focal point, with new datasets and benchmarks being established to test models against a wider range of synthetic audio scenarios (Yi et al., 2022; Müller et al., 2022).
2.2 Continual Learning Methods
Continual learning, also known as lifelong learning, addresses the challenge of training models on sequential tasks while retaining knowledge from previous tasks. In the context of deepfake audio detection, this capability is crucial as new types of attacks and spoofing techniques emerge. Traditional models suffer from catastrophic forgetting, where learning new tasks leads to the erosion of previously acquired knowledge. To mitigate this, methods such as Elastic Weight Consolidation (EWC) have been proposed, which constrain the updates of important parameters to preserve knowledge from earlier tasks (Kirkpatrick et al., 2017). Other techniques, such as experience replay (Rolnick et al., 2019) and knowledge distillation (Li & Hoiem, 2017), have also been employed to maintain performance across tasks by rehearsing previous data or transferring knowledge from a stable teacher model to a student model during training. Recent advancements in continual learning have focused on adapting these methods to complex neural architectures, such as transformers and deep CNNs, making them more robust in dynamic environments where the nature of the data and tasks evolves over time (Rebuffi et al., 2017). In the context of deepfake detection, these continual learning approaches (Zhang et al., 2023; 2024) are pivotal in ensuring that models can adapt to new challenges without (or miner) losing the ability to detect previously known spoofing techniques.
3 Proposed Method
To address the growing complexity and diversity of audio deepfake detection, we propose a comprehensive benchmark designed to rigorously evaluate the performance of anti-spoofing models across a variety of challenging scenarios. Existing audio deepfake detection methods often focus on limited conditions or datasets, failing to capture the full range of real-world variations that models may encounter (Ma et al., 2021; Zhang et al., 2023; 2024). Our benchmark aims to fill this gap by incorporating eight distinct tasks, each carefully selected to represent different data sources, languages, and acoustic conditions. This diversity allows for a thorough assessment of model robustness, adaptability, and cross-lingual capabilities, pushing the boundaries of current audio spoofing detection methods. The benchmark is structured to challenge models with varying degrees of difficulty, ensuring that only the most advanced and resilient models can succeed across all tasks.
3.1 Data Architecture
Our benchmark architecture consists of eight tasks (Task 1 to Task 8), each representing different datasets and conditions for detecting synthetic audio. These tasks vary in data sources, language, acoustic conditions, and keep the number of training and evaluation samples in same. The data architecture of EVDA benchmark are listed as Table 1, which is designed to comprehensively evaluate the performance of audio spoofing and anti-spoofing models across a diverse set of conditions and datasets. Below, we detail each experiment’s setup and objectives:
Table 1: Data architecture of our proposed benchmark
ID Source Have Vocoder ID Language Acoustic Condition Train Test
Task 1 FMFCC-A F Chinese Human 2000 5000
Task 2 In-the-Wild F English Real world 2000 5000
Task 3 ADD 2022 F Chinese Low-quality/Partially fake 2000 5000
Task 4 ASVspoof2015 T English No significant noise 2000 5000
Task 5 ASVspoof2019LA T English Logical access 2000 5000
Task 6 ASVspoof2021LA T English Logical access 2000 5000
Task 7 FoR F English Human 2000 5000
Task 8 HAD F Chinese Human/Partially fake 2000 5000
Task 1: FMFCC-A (Chinese)(Zhang et al., 2021) - Utilizes 2000 training and 5000 evaluation samples of Chinese audio. This experiment is critical for models to adapt to low-quality, partially fake audio environments, increasing the difficulty of the spoofing detection task.
Task 2: In-the-Wild (English)(Müller et al., 2022) - Uses 2000 training and 5000 evaluation samples in real-world audio scenarios, testing the models’ effectiveness in dynamically changing and challenging acoustic environments.
Task 3: ADD 2022 (Chinese)(Yi et al., 2022) - Incorporates 2000 training and 5000 evaluation samples of Chinese audio. This experiment focuses on low-quality, partially fake settings, presenting challenges in detecting spoofed audio.
Task 4: ASVspoof2015 (English)(Wu et al., 2015) - Utilizes 2000 training and 5000 evaluation samples from the ASVspoof2015 dataset, focusing on English language audio in clean conditions without significant channel or background noise effects. Vocoder ID is available, aiding in more precise synthetic speech detection.
Task 5: ASVspoof2019 - Logical Access (LA) (English)(Todisco et al., 2019) - Incorporates 2000 training and 5000 evaluation samples focusing on logical access scenarios to test models against sophisticated spoofing attacks in English. Vocoder ID is again available, providing critical data for assessing synthetic signal traits.
Task 6: ASVspoof2021 - Logical Access (LA) (English)(Liu et al., 2023) - Continues the logical access examination with the latest dataset iteration, using English audio samples. The presence of vocoder ID helps in identifying new spoofing techniques that have emerged since the 2019 dataset.
Task 7: FoR (English)(Reimao & Tzerpos, 2019) - Uses 2000 training and 5000 evaluation samples of purely human-generated audio in English, testing the robustness of models against non-synthetic audio scenarios.
Task 8: HAD (Chinese)(Yi et al., 2021) - Utilizes 2000 training and 5000 evaluation samples of Chinese audio, which includes a mix of human and partially fake audio, presenting challenges in cross-lingual and authenticity discernment.
3.2 Arrangement of the tasks
We have arranged and highlighted the table data based on the distinct characteristics of each experiment on Table 1, emphasizing the differences in language and acoustic conditions. This organization allows for a clearer comparison of how various datasets and conditions affect the performance of deepfake audio detection models. By highlighting the changes, we can better understand the specific challenges each dataset introduces and how the models need to adapt to these variations. The key distinctions in the experiments are reflected in the changes to both language and acoustic conditions as follows:
- Task 1 to Task 2: The language changes from Chinese to English, with the acoustic condition shifting from ”Human” to ”Real world.” This transition highlights the model’s need to adapt from a controlled environment to more natural, varied conditions.
- Task 2 to Task 3: The language switches back to Chinese, while the acoustic condition changes to ”Low-quality/Partially fake,” indicating a mix of real and synthetic audio in more challenging quality conditions.
- Task 3 to Task 4: The language changes to English, with the acoustic condition becoming ”No significant noise,” focusing on clean audio without significant interference.
- Task 4 to Task 5: The language remains English, but the focus shifts to ”Logical access” scenarios, introducing sophisticated spoofing attacks.
- Task 5 to Task 6: The language and acoustic condition remain consistent (English, ”Logical access”), but the dataset iteration updates to the latest version with newly emerged spoofing technologies.
- Task 6 to Task 7: The language stays the same (English), while the acoustic condition returns to ”Human,” emphasizing purely human-generated audio.
- Task 7 to Task 8: The language changes back to Chinese, with the acoustic condition now being ”Human/Partially fake,” involving a mix of human and synthetic audio.
These highlighted changes are crucial for understanding the adaptability and robustness of deepfake audio detection models across different datasets, languages, and acoustic conditions. The benchmark is designed not only to assess the current state-of-the-art in spoofing detection but also to push the development of more robust, adaptable, and cross-lingual models capable of handling the complexities of real-world audio spoofing scenarios.
3.3 Data Selection Methodology
For each task within our benchmark, we provided two strategies for data selection:
Random Selection: We randomly select a fixed number of authentic and spoofed audio samples from the original datasets or collected data. Initially, we split the data into two groups: Real and Fake. From the ”Real” group, we randomly select 1000 real samples, and from the ”Fake” group, we select 1000 fake samples corresponding to various generation methods. These samples are combined to form the training set. For the evaluation set, we follow the same procedure but select only 2500 real and 2500 fake samples. This approach ensures a balanced representation of real and fake samples in both the training and evaluation sets. In current benchmark, this method has been used for data selection.
Most Informative Selection: In this method, we aim to select the most informative samples from the original datasets. First, we train $n$ discriminative models for spoof detection using external datasets, ensuring these models achieve extremely low loss on their respective training sets. These well-trained models act as experts during the sample selection process. Each sample from the original dataset is evaluated by all experts, and the samples are scored based on the entropy of the experts’ predictions using the formula:
$\text{Entropy}=-(p_{1}\log p_{1}+p_{2}\log p_{2})$ (1)
where
$p_{1}=\frac{N_{\text{real}}}{N_{\text{real}}+N_{\text{fake}}},\quad p_{2}=\frac{N_{\text{fake}}}{N_{\text{real}}+N_{\text{fake}}}$ (2)
Here, $N_{\text{real}}$ and $N_{\text{fake}}$ represent the number of models classifying the sample as real or fake, respectively. According to information theory (Shannon, 1948), samples with the highest entropy values contain the most information, indicating uncertainty and hence, difficulty in classification. We list the entropy of each sample and rank them from highest to lowest. The top 2500 real and 2500 fake samples are selected to construct the evaluation set, ensuring it includes the most challenging and informative samples. The next 1000 real and 1000 fake samples are selected for the training set, which benefits from the inclusion of diverse and informative samples.
3.4 Continuous Learning Methods in proposed Benchmark
The benchmark evaluates various continual learning methods to enhance the adaptability of audio spoofing detection systems. These approaches are summarized as follows:
Replay (Rolnick et al., 2019) mitigates catastrophic forgetting by replaying a subset of previously seen samples during training, helping to maintain performance on earlier tasks.
Finetuning adapts a pre-trained model to new tasks by continuing training with new data, effectively leveraging existing models to incorporate new knowledge.
EWC (Kirkpatrick et al., 2017) adds a regularization penalty to the loss function based on parameter importance, preserving critical knowledge from previous tasks.
GDumb (Prabhu et al., 2020) combines finetuning with periodic retraining on a data subset, balancing new and old information for large-scale learning.
CWRStar (Maltoni & Lomonaco, 2019) employs rehearsal and interleaving tasks using a fixed memory, improving retention of past knowledge while learning new tasks.
SI (Zenke et al., 2017) introduces dynamic regularization based on the importance of model parameters, balancing the retention of old knowledge with the acquisition of new information.
OWM (Zeng et al., 2019) adjusts weight updates to remain orthogonal to those associated with previous tasks, reducing forgetting by preserving relevant features.
RAWM (Zhang et al., 2023) adapts weight updates based on the ratio of genuine to fake utterances, introducing regularization to retain knowledge across different acoustic conditions.
RWM (Zhang et al., 2024) uses radian-based weight modification to minimize disruption to previously learned features, enhancing stability and adaptability in continual learning.
In summary, our benchmark provides a comprehensive evaluation of these methods, aiming to develop robust models that can effectively handle evolving audio spoofing attacks with high reliability and accuracy.
4 Result
4.1 Compared with multiply continual learning methods
Methods Task1 Task2 Task3 Task4 Task5 Task6 Task7 Task8 Avg
Replay 2.88 7.60 2.56 1.40 2.80 8.84 6.76 4.12 4.62
Finetuning 8.28 10.04 5.36 1.52 4.08 18.24 10.44 4.44 7.80
EWC 6.64 10.16 5.52 1.72 4.20 16.92 10.08 4.56 7.475
GDumbFinetune 2.88 8.28 4.92 3.64 5.88 19.28 13.16 10.44 8.56
CWRStar 28.2 36.72 22.48 29.8 46.4 49.36 47.04 34.12 36.77
SI 8.52 10.2 7.48 1.68 5.40 17.12 13.92 4.60 8.615
OWM 6.72 28.52 32.88 14.32 19.84 40.8 26.08 21.88 23.88
RAWM 9.24 13.6 5.38 1.68 4.36 16.4 12.44 6.00 8.638
RWM 9.64 14.76 8.52 2.04 4.64 17.6 11.88 5.96 9.38
Table 2: EER (%) on each task for different continual learning methods. The cells with the lowest EER in each column are highlighted in red, while the cells with the second-lowest EER are highlighted in yellow. Lower EER values indicate better performance.
Table 2 summarizes the performance of various continual learning methods across eight tasks, evaluated by Equal Error Rate (EER) (Wu et al., 2015). Compared with all continual learning methods, replay consistently shows most competitive performance across all tasks, demonstrating its overall effectiveness. Finetuning and EWC exhibit relatively similar performances. However, EWC achieves the better performance in old and average tasks, indicating its overcoming forgetting performance in old tasks. GDumbFinetune performs well in certain tasks, notably achieving the lowest EER in Task 1 along with Replay. However, its overall performance is less consistent, with a higher average EER. CWRStar consistently shows higher EER values across all tasks, indicating it is less effective compared to other methods, with the highest average EER (36.77).SI and OWM demonstrate mixed results. SI achieves low EERs in some tasks, but not as consistently as Replay or EWC. OWM shows the highest EER in several tasks, particularly in Task2 and Task3. RAWM and RWM provide relatively competitive performances but fall behind Replay and EWC in several tasks. RAWM has a notable second-lowest EER in Task 6 compared to the best. Overall, Replay and EWC are the top-performing methods, with Replay showing the lowest average EER. These results suggest that Replay and EWC are effective in minimizing error rates in continual learning scenarios for audio deepfake detection, while other methods like CWRStar and OWM may need further refinement for better performance.
4.2 Continual learning on each task
We evaluate our benchmark using the EER metric on 8 tasks using the conventional Finetuning and EWC continual learning method. All methods and tasks are trained using 5 linear layers model with 128-hidden dimension. The EER results are illustrated as Table 3:
Train after Task1 Task2 Task3 Task4 Task5 Task6 Task7 Task8
Task1 FT 5.44 32.12 28.52 33.80 42.12 39.20 45.00 22.32
EWC 5.44 32.12 28.52 33.80 42.12 39.20 45.00 22.32
Task2 FT 10.20 6.32 25.84 7.72 10.64 30.60 21.92 12.28
EWC 9.48 6.20 26.72 8.96 12.24 29.92 23.00 12.24
Task3 FT 11.48 15.04 1.48 19.16 44.40 34.92 46.28 19.72
EWC 12.92 15.36 1.64 20.40 49.80 36.84 42.76 20.16
Task4 FT 16.56 13.00 4.12 1.80 6.32 25.76 17.36 9.92
EWC 17.96 12.80 4.80 1.72 6.16 27.32 16.76 9.92
Task5 FT 18.56 10.28 9.28 1.76 2.64 23.60 12.08 11.08
EWC 21.04 9.96 10.60 1.68 2.36 23.88 12.16 10.96
Task6 FT 10.04 12.36 5.64 2.40 6.96 9.96 14.12 11.16
EWC 10.12 11.52 6.20 2.16 6.52 10.04 13.84 11.52
Task7 FT 10.84 12.72 8.20 1.72 2.64 12.44 6.64 12.40
EWC 11.08 11.96 9.72 2.04 2.52 12.84 6.28 12.64
Task8 FT 8.28 10.04 5.36 1.52 4.08 18.24 10.44 4.44
EWC 6.64 10.16 5.52 1.72 4.20 16.92 10.08 4.56
Table 3: The EER(%) evaluated on all tasks after each training using conventional finetuning (FT) and EWC.
The results presented in Table 3 compare the EER (%) across all tasks after each training session using two methods: conventional finetuning (FT) and Elastic Weight Consolidation (EWC). From the table, it is evident that both methods exhibit similar performance when applied to the first task, as expected, since no prior knowledge needs to be preserved. However, as training progresses on subsequent tasks, noticeable differences emerge between the two methods. For Task 2, EWC shows slightly improved performance over FT in both Task 1 and Task 2, with reductions in EER ranging from minor differences. These improvements suggest that EWC effectively mitigates catastrophic forgetting by preserving knowledge from earlier tasks, allowing for more stable learning. However, there are cases, like Task 3, where EWC does not offer significant improvements over FT, indicating that its effectiveness can be task-dependent. Overall, while both methods show varying levels of performance across different tasks, EWC generally provides a modest improvement over FT on old tasks. This highlights advantage in continual learning where the model needs to maintain performance across multiple tasks without revisiting them.
5 Discussion
In this work, we have proposed a continual learning benchmark, which is the first as best as we know, for audio deepfake detection. One of the key advantages of our work is the thorough evaluation of continual learning methods in a challenging domain-shift condition. The varied and complex nature of audio data, combined with the evolving techniques used in generating deepfakes, makes it a particularly demanding area for continual learning. Our results provide valuable insights into the current capabilities of these methods and their applicability to real-world problems.
However, our study also has certain limitations. Firstly, the scope of continual learning methods explored in this work is limited. Although we have focused on well-established methods like FT and EWC, there are several other approaches that could offer additional benefits, such as replay-based methods, regularization techniques, and meta-learnin approaches. Furthermore, the benchmark tasks utilized in this study, while representative, may not fully capture the diversity of audio deepfakes encountered in practice.
In future work, we plan to expand our investigation to include a broader range of continual learning methods. This will not only provide a more complete understanding of the strengths and weaknesses of different approaches but also enable the proposal of novel methods tailored specifically to the challenges of audio deepfake detection. Additionally, we intend to incorporate newly emerging audio deepfake techniques, such as those powered by advanced generative models like GPT-4 and its successors, into our evaluation framework. By doing so, we aim to ensure that our models remain robust and effective in detecting the latest deepfake techniques, thereby maintaining their relevance in the face of rapidly advancing technologies.
6 Conclusion
In conclusion, the rapid evolution of audio generation models poses significant challenges for fake audio detection. However, by adopting continual learning approaches and establishing a comprehensive benchmark, we can effectively address these challenges. Our proposed benchmark provides a standardized platform for evaluating detection methods against evolving synthetic audio techniques, facilitating the development of more robust and effective detection algorithms.